# A Meat-Summer Night’s Dream: A Tangible Design Fiction Exploration of Eating Biohybrid Flying Robots

### v1

@@ Abstract.
What if future dining involved eating robots? We explore this question through a playful and poetic experiential dinner theater: a tangible design fiction staged as a 2052 Paris restaurant where diners consume a biohybrid flying robot in place of the banned delicacy of ortolan bunting. Moving beyond textual or visual speculation, our “dinner-in-the-drama” combined performance, ritual, and multisensory immersion to provoke reflection on sustainability, ethics, and cultural identity. Six participants from creative industries engaged as diners and role-players, responding with curiosity, discomfort, and philosophical debate. They imagined biohybrids as both plausible and unsettling—raising questions of sentience, symbolism, and technology adoption that exceed conventional sustainability framings of synthetic meat. Our contributions to HCI are threefold: (i) a speculative artifact that stages robots as food, (ii) empirical insights into how publics negotiate cultural and ethical boundaries in post-natural eating, and (iii) a methodological advance in embodied, multisensory design fiction.
@@
copyright: rightsretainedjournalyear: 2025ccs: Human-centered computing Empirical studies in HCI
Figure 1. Participants were eating “biohybrid flying robots” in the traditional French manner.:
@@ 1. Introduction
— Jules Verne, French writer, “father of science fiction”.
@@
— Marvin Minsky, American cognitive and computer scientist, AI pioneer.
@@
Since time immemorial, humans have eaten animals. This act—mundane yet deeply cultural—has shaped our ethics, our rituals, and our relationship to the living world. Today, however, climate change, industrial farming, and shifting cultural values press us to imagine new food futures. Increasingly, researchers are exploring radically alternative possibilities. For example, the RoboFood project (RoboFood, nd) exemplifies a broader movement where robotics and food science converge around the idea of edible robotics to reimagine what counts as food. Set against such innovations and imaginaries we ask: what if we could eat biohybrid robots—robots that are (partly) made from cultured tissues—instead of cows, chickens, or fish? In this paper, we take this speculative question constructively, not to propose a technological roadmap, but to use it as a lens: a way to probe how emerging technologies might reshape human ethics, cultural identity, and the boundaries of human–robot interaction.
@@
Although eating animals has been a constant throughout evolutionary epochs, successive sociotechnical revolutions have transformed this relationship. Technologies have taken us from hunting “wild” animals to domestication, selective breeding (Norwood and Lusk, 2011), and eventually mass production. More recently, an emergent technology of cultivated or lab-grown meat has reached commercial approval. At present, Singapore and the USA are the only two countries that have legalized the sale of lab-grown meat for human consumption, with others likely to follow. For instance, the UK has approved lab-grown meat for pet food (Speed, 2024), and the Netherlands permits it for tasting purposes (Verhulst, 2023). These developments mark the beginning of a food revolution (Califf and Mayne, 2022), one that seeks to address the environmental and ethical costs of industrial meat production.
@@
Yet the state of the art remains at an early stage, with technical, socio-political, and regulatory challenges ahead (Stephens et al., 2018). A key difficulty lies in replicating the form, texture, and appearance of conventional animal meat (Raman, 2021), given the complexity of biological structures (Griffith and Naughton, 2002). This leads us to ask: Why not consider creating a biohybrid robot that leverages natural processes to grow authentic animal components, while being managed and controlled by computers? Here, Minsky’s metaphor of the brain as a “meat machine” offers a useful conceptual pivot. By framing the brain—the seat of consciousness—as mechanistically reproducible, Minsky underscored a materialist view that, in principle, no natural phenomenon is beyond technological replication. Extending this reasoning, one might imagine not only replicating neural systems but also reconstituting organic tissues, organs, and even entire animal forms within artificial substrates. This conceptual foundation supports the possibility of assembling animal–robot hybrids: entities that merge biological growth with computational control. From this perspective, biohybrid robots are not merely a technical workaround for the shortcomings of lab-grown meat. They represent a deeper continuity with cultural traditions of animals-as-food, while simultaneously reframing the act of eating within the context of human–robot interaction.
To bring this vision to life, we designed an experiential dinner theater titled A Meat-Summer Night’s Dream. Inspired by Shakespeare’s A Midsummer Night’s Dream, the event blended performance, fine dining, and speculative narrative to critically explore future food rituals and revive a fading culinary tradition through speculative robotics and bioengineering. Specifically, we imagined a future where synthetic meat, grown as part of biohybrid flying robots, replaces the controversial French delicacy, Ortolan Bunting. This multisensory performance was not intended to propose a product or policy, but to provoke critical reflection and open dialogue around food ethics, technological futures, and cultural preservation.
Situated in Paris in 2052, the “dinner-in-the-drama” invited six participants who were not technical experts, but from the creative industries to embody roles within a futuristic gastronomic ritual in an art gallery setting. This tangible design fiction provided an embodied context to examine how people engage with speculative technologies when immersed in rich cultural and sensory environments. To guide our inquiry, we posed two research questions: RQ1: How do people contemplate the idea of eating biohybrid robots when exploring alternative food futures? RQ2: How does a tangible multisensory dining experience provoke critical reflection around sustainable food futures?
This paper contributes to HCI: (i) by presenting a speculative artifact—an envisionment and partial enactment of consuming biohybrid robots in future food rituals—that provokes critical reflection and discussion around alternative food futures and their ethical and sustainability implications; (ii) by offering empirical insights from participants’ responses to a multisensory, performative experience, which surfaced reflections on sentience, sustainability, and cultural meaning; and (iii) through methodological development, by adapting design fiction into a tangible, embodied format that fosters critical, experiential engagement with speculative technologies.
@@ 2.1.2. Synthetic Food: Cultivated Meat and Its Limitations
Despite this promise, cultivated meat faces significant scientific and socio-political barriers. Key technical limitations include replicating the complex, hierarchical structures of natural tissues—integrating microscale functional units into macroscale edible forms remains an ongoing challenge (Griffith and Naughton, 2002). Furthermore, regulatory, cultural, and consumer acceptance issues persist (Stephens et al., 2018).
@@ 2.2. The Idea of Living Machines, Biohybrid Robots, and the Animal-Robot Analogy
Over the last decade, the term Living Machines has been employed to signal a growing convergence between biology and technology (see (Prescott et al., 2014, 2018)). In both scientific and artistic domains, researchers have developed machines inspired by biological forms and processes (Wang et al., 2023b, 2022; Karásek et al., 2018; Wang et al., 2025a), and increasingly, machines built from living tissue itself (Morimoto et al., 2018; Kawai et al., 2024; Kinjo et al., 2024). For instance, Takeuchi and colleagues created a biohybrid bipedal robot powered by skeletal muscle tissue (Kinjo et al., 2024) and a humanoid robot covered with living skin capable of smiling (Kawai et al., 2024). Other works employ fungi as computational substrates for robotic control (Mishra et al., 2024), blurring the boundaries between machine and organism. At the same time, ethical challenges and responsibilities in biohybrid robotics research have been raised (Mestre et al., 2024; Kirksey, 2021), as these efforts increasingly collapse distinctions between the biological and the mechanical. These efforts are part of a broader movement to develop biohybrid systems, where biological and synthetic components form new composite entities (Jung, 2011; Webster-Wood et al., 2022). Within this movement, edible robotic concepts (Floreano et al., 2024) also extend the scope of “living” or “nutritive” machines, challenging the division between machine, organism, and food.
@@ 3.2. Dinner-in-the-Drama as Tangible Design Fiction
Our approach is motivated by the observation that human experience, especial culinary practices, is inherently multisensory: touch, smell, sound, and taste shape how we perceive and emotionally respond to the world. By engaging these sensory modalities, tangible design fiction narrows the distance between speculative imagination and physical experience, enabling participants to feel, not just imagine, a possible futures.
@@
3.3. Drama Design and Prototyping
The visual identity of the event was themed around the color blue, symbolizing sky, freedom, and artificiality. This motif extended from the lighting and interior setup to drinks and desserts, invoking a sense of alien futurity. Table settings included napkins (for the Ortolan ritual), feathers, bird dolls, and glassware shaped like birds. These theatrical elements were chosen for their symbolic and sensory resonance, blending natural motifs with synthetic speculation. All guests were instructed to place napkins over their heads during the main course, reviving the traditional Ortolan ritual in a new technological context.
Figure 2. Theatrical props showcasing the biohybrid flying robot in three different stages: (a) Flying drone (in motion); (b) 3D-printed prototype; (c) Roasted bird (edible).:
Culinary Design and Symbolism: The three-course menu (see Appendix A) reflected the narrative arc of the drama and incorporated speculative and symbolic ingredients. The starter mimicked the natural diet of Ortolans, placing participants in the imagined perspective of the bird. The main course substituted Ortolan with quail for ethical, legal, and sensory alignment, allowing diners to experience a taste analogous to the original dish (Poon, 2014). The final dessert—a futuristic blue concoction—reinforced the speculative setting and closed the meal with a sense of surrealism. Real ingredients were disclosed in advance to respect dietary needs and ethical transparency.
3.4. Participants and Their Role Play
We recruited six guest participants using snowball sampling, all professionals working in creative, design, and cultural sectors. This selection was intentional: individuals with imaginative capacities and critical reflection skills were likely to engage productively with the speculative scenario. The participants were: P1, a design lecturer at a world-renowned institution; P2, an award-winning poet; P3, an artist and gallery owner; P4, a foresight expert from a global tech innovation hub; P5, a professional UX designer; and P6, a filmmaker. The group was gender-balanced, with five aged 30s and one aged 70s. Participants were verified as meat eaters, and food allergies were checked in advance. One participant with dietary restrictions participated as a non-eater, taking on the role of the restaurant service manager.
@@ 3.5. Event Procedure
Figure 3. Procedure diagram.:
@@
The lead artist (host) welcomed guests and introduced the speculative narrative. The chef (researcher) then explained the fictional menu and contextualized the key concept of edible animal-robot hybrids. Dramatic beats followed: a feathered drone representing the “synthetic ortolan” was flown across the room (Figure 4d), after which it was replaced by a 3D prototype passed among the guests (Figure 4e). The final edible course—a roasted quail representing the hybrid—was served under napkins, referencing the historic Ortolan tradition (Figure 4f). The experience concluded with a dessert and a moderated post-dinner discussion.
Figure 4. Highlights from the dinner-in-the-drama: (a) Opening remarks; (b) Introducing key concepts; (c) Enjoying the show; (d) Biohybrid robot in flight; (e) Observing the biohybrid robot; (f) Serving biohybrid robots; (g) Eating biohybrid robots.:

### v2

@@ Abstract.
What if future dining involved eating robots? We explore this question through a playful and poetic experiential dinner theater: a tangible design fiction staged as a 2052 Paris restaurant where diners consume a biohybrid flying robot in place of the banned delicacy of ortolan bunting. Moving beyond textual or visual speculation, our “dinner-in-the-drama” combined performance, ritual, and multisensory immersion to provoke reflection on sustainability, ethics, and cultural identity. Six participants from creative industries engaged as diners and role-players, responding with curiosity, discomfort, and philosophical debate. They imagined biohybrids as both plausible and unsettling—raising questions of sentience, symbolism, and technology adoption that extend beyond conventional sustainability framings of synthetic meat. Our contributions to HCI are threefold: (i) a speculative artifact that stages robots as food, (ii) empirical insights into how people negotiate cultural and ethical boundaries in post-natural eating, and (iii) a methodological advance in embodied, multisensory design fiction.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790678isbn: 979-8-4007-2278-3/2026/04ccs: General and reference Designccs: Human-centered computing Empirical studies in HCIccs: Computer systems organization Roboticsccs: Applied computing Performing artsccs: Hardware Emerging interfaces
Figure 1. Participants were eating “biohybrid flying robots” in the traditional French manner.
A group of people eating at a table with napkins covering their heads.
@@ 1. Introduction
— Jules Verne,
French writer,
“father of science fiction”.
@@
— Marvin Minsky,
American cognitive and computer scientist,
AI pioneer.
@@
Since time immemorial, humans have eaten animals. This act—mundane yet deeply cultural—has shaped our ethics, our rituals, and our relationship to the living world. Today, however, climate change, industrial farming, and shifting cultural values press us to imagine new food futures. Increasingly, researchers are exploring radically alternative possibilities. For example, the ROBOFOOD project (RoboFood, n.d.) exemplifies a broader movement where robotics and food science converge around the idea of edible robotics to reimagine what counts as food. These shifts create new ethical and cultural tensions that invite HCI to interrogate how technological interventions might reshape long-standing food rituals and the meaning of “eating” itself. We ask: what if we could eat biohybrid robots—robots that are (partly) made from cultured tissues—instead of cows, chickens, or fish? In this paper, we take this speculative question constructively, not to propose a technological roadmap, but to use it as a lens: a way to probe how emerging technologies might reshape human ethics, cultural identity, and the boundaries of human–robot interaction.
@@
Although eating animals has been a constant throughout evolutionary epochs, successive sociotechnical revolutions have transformed this relationship. Technologies have taken us from hunting wild animals to domestication, selective breeding (Norwood and Lusk, 2011), and eventually mass production. More recently, an emergent technology of cultivated or lab-grown meat has reached commercial approval. At present, Singapore and the USA are the only two countries that have legalized the sale of lab-grown meat for human consumption, with others likely to follow. For instance, the UK has approved lab-grown meat for pet food (Speed, 2024), and the Netherlands permits it for tasting purposes (Verhulst, 2023). These developments mark the beginning of a food revolution (Califf and Mayne, 2022), one that seeks to address the environmental and ethical costs of industrial meat production.
@@
Yet the state of the art remains at an early stage, with technical, socio-political, and regulatory challenges ahead (Stephens et al., 2018). A key difficulty lies in replicating the form, texture, and appearance of conventional animal meat (Raman, 2021), given the complexity of biological structures (Griffith and Naughton, 2002). This leads us to ask: Why not consider creating a biohybrid robot that leverages natural processes to grow authentic animal components, while being managed and controlled by computers? Here, Minsky’s metaphor of the brain as a “meat machine” offers a useful conceptual pivot. By framing the brain—the seat of consciousness—as mechanistically reproducible, Minsky underscored a materialist view that, in principle, no natural phenomenon is beyond technological replication. Extending this reasoning, one might imagine not only replicating neural systems but also reconstituting organic tissues, organs, and even entire animal forms within artificial substrates. This conceptual foundation supports the possibility of assembling and eating animal–robot hybrids (Wang et al., 2023a): entities that merge biological growth with computational control. From this perspective, such biohybrid robots are not merely a technical workaround for the shortcomings of current lab-grown meat. They represent a deeper continuity with cultural traditions of animals-as-food, while simultaneously reframing the act of eating within the context of human–robot interaction.
To bring this vision to life, we designed an experiential dinner theater titled A Meat-Summer Night’s Dream, set in a fictional 2052 Paris restaurant. Six participants from creative and cultural fields engaged as diners and semi-scripted role-players as they encountered a series of tangible speculative artifacts—including a feathered drone, a translucent anatomical model, and an edible surrogate bird (see Figure 2). Unlike conventional design fictions that rely on texts or images, our approach foregrounds multisensory experience—taste, smell, touch, sound, and visual occlusion—to elicit situated ethical reflection. We additionally foreground the cultural stakes of eating by reimagining the controversial French delicacy of Ortolan Bunting, substituting the bird with synthetic meat grown as part of biohybrid flying robots. This choice is not celebratory; rather, it surfaces questions of ritual, taboo, symbolism, and cultural identity that are often under-examined in HCI discussions of food futures. To guide our inquiry, we posed two research questions: RQ1: How do people contemplate the idea of eating biohybrid robots when placed within a culturally and sensorially rich speculative future? RQ2: How does a tangible multisensory dining experience provoke ethical, cultural, and affective reflection about post-natural food futures?
This paper contributes to HCI in three ways: (i) a situated, embodied speculative artifact that employs sensory, material, and ritual elements to explore possible interactions with edible biohybrid robots; (ii) empirical finding that illuminate how participants reason about animacy, ethical ambiguity, and cultural meaning within a multisensory speculative encounter; and (iii) a methodological refinement of design fiction, demonstrating how a tangible, embodied format can foster critical, experiential engagement with speculative technologies.
Finally, we clarify the scope of the work—this study does not aim for generalizability or predictive claims. Rather, its value lies in documenting how speculative dining can elicit situated ethical and cultural reasoning in an intimate, embodied context. We recognize that this paper’s provocative nature has sparked and may continue to spark diverse reactions, and we deeply appreciate this range of perspectives. We argue that such plurality is itself evidence of the value of speculative and experimental work in HCI, namely its capacity to open questions, stimulate discussion, and broaden our imagination of what interaction can be.
@@ 2.1.2. Synthetic Food: Cultivated Meat and Its Limitations
Despite this promise, cultivated meat faces significant scientific and socio-political barriers. Key technical limitations include replicating the complex, hierarchical structures of natural tissues, as integrating microscale functional units into macroscale edible forms remains an ongoing challenge (Griffith and Naughton, 2002). Furthermore, regulatory, cultural, and consumer acceptance issues persist (Stephens et al., 2018).
@@ 2.2. The Idea of Living Machines, Biohybrid Robots, and the Animal-Robot Analogy
Over the last decade, the term Living Machines has been employed to signal a growing convergence between biology and technology (see (Prescott et al., 2014, 2018)). In both scientific and artistic domains, researchers have developed machines inspired by biological forms and processes (Wang et al., 2023b, 2022; Karásek et al., 2018; Wang et al., 2025a, 2024; Wang, 2025), and increasingly, machines built from living tissue itself (Morimoto et al., 2018; Kawai et al., 2024; Kinjo et al., 2024). For instance, Takeuchi and colleagues created a biohybrid bipedal robot powered by skeletal muscle tissue (Kinjo et al., 2024) and a humanoid robot covered with living skin capable of smiling (Kawai et al., 2024). Other works employ fungi as computational substrates for robotic control (Mishra et al., 2024), blurring the boundaries between machine and organism. At the same time, ethical challenges and responsibilities in biohybrid robotics research have been raised (Mestre et al., 2024; Kirksey, 2021), as these efforts increasingly collapse distinctions between the biological and the mechanical. These efforts are part of a broader movement to develop biohybrid systems, where biological and synthetic components form new composite entities (Jung, 2011; Webster-Wood et al., 2022). Within this movement, edible robotic concepts (Floreano et al., 2024) also extend the scope of “living” or “nutritive” machines, challenging the division between machine, organism, and food.
@@ 2.3. Speculative Design and Design Fictions in HCI
2.4. Cultures of Edibility and Biohybrid Ethics
Researchers have long argued that food ethics and cultural practices are co-constitutive: what counts as acceptable eating is shaped by collective identities, social memories, and moral orders (Fischler, 1988; Counihan, 2013; Douglas, 1972). Synthetic and biohybrid foods enter these domains not as neutral technologies but as culturally interpreted matter entangled with histories of purity, taboo, kinship, and ritual (Rozin et al., 2008; Harris, 1987, 2001). Studies in food anthropology show that unfamiliar foods are typically assimilated through analogy, symbolic categorization, and everyday narrative rather than technical knowledge (Sutton, 2006; PAXSON, 2008; Mintz, 1986), indicating that interpretation is grounded in accumulated cultural repertoires and affective attachments.
Affective and sensory dimensions are central to this negotiation. Previous works (Probyn, 2016; Highmore, 2008) show that disgust, curiosity, care, and desire stabilize decisions about edibility. Multispecies ethics research further argues that judgments about consuming hybrid organisms hinge on perceived agency, liveliness, and relationality (Haraway, 2008; Braidotti, 2013). The introduction of semi-living tissues and biohybrid robotics complicates conventional distinctions between life and mechanism, raising questions familiar from stem-cell ethics, embryology, and synthetic biology (Hyun et al., 2016; for Stem Cell Research, 2021; Webster-Wood et al., 2022).
STS scholarship highlights that ethical reasoning around food technologies is embedded in broader political and ecological imaginaries. Analyses of food futures emphasize how visions of climate adaptation, industrial scaling, and geopolitical distribution shape acceptance of synthetic and lab-grown products (Gobbo, 2023; Ayakannu et al., 2024; tamás_landesz_2023). Paxson’s post-pasteurian microbiopolitics (PAXSON, 2008) shows that novel food systems produce new forms of governance, expertise, and public trust. Star and Griesemer’s concept of boundary objects (Star and Griesemer, 1989) has been applied to food technologies as hybrid entities negotiated across scientific, culinary, and moral communities (Bingham, 2006), framing edibility as a provisional agreement rather than an intrinsic property.
Recent speculative and sensory design research provides methodological insight into how publics encounter ethical ambiguities before technologies are widely implemented. Immersive and anticipatory design methods reveal latent values and cultural scripts that remain obscure in abstract discussion (Auger, 2013; Michael, 2012; Jewitt et al., 2022). Lupton and Willis (Lupton and Willis, 2021; Lupton, 2022) demonstrate that participatory, embodied encounters offer “anticipatory ethics”, allowing participants to rehearse future dilemmas using familiar cultural frameworks. Related multisensory design work shows that taste, smell, and ritual materially shape ethical reasoning, not merely represent it (Pink, 2023).
Existing scholarship makes clear that judgments of edibility emerge through cultural infrastructures, multispecies ethics, and affective reasoning, yet these dynamics are mostly examined retrospectively. What remains less understood is how publics might actively negotiate ethical ambiguities before biohybrid foods become commonplace. Our study addresses this gap by using speculative dining to elicit situated ethical reasoning through ritual, narrative, and sensory engagement, positioning culture as an active arena where meanings of biohybridity and synthetic eating are negotiated in real time.
@@ 3.1.2. Biohybrid Flying Robots as Synthetic Ortolan for Eating
From left to right: a flying robot with blue feather, a colorful resin prototype, a bird dish with fruits and vegetables on a plate.
Figure 2. Theatrical props showcasing the biohybrid flying robot in three different stages: (a) Flying drone (in motion); (b) 3D-printed prototype; (c) Roasted bird (edible).
@@ 3.2. Dinner-in-the-Drama as Tangible Design Fiction
Our approach is motivated by the observation that human experience, especially culinary practices, is inherently multisensory: touch, smell, sound, and taste shape how we perceive and emotionally respond to the world. By engaging these sensory modalities, tangible design fiction narrows the distance between speculative imagination and physical experience, enabling participants to feel, not just imagine, a possible future.
@@
3.3. Drama Design and Prototyping
The visual identity of the event was themed around the color blue, for its associations with the sky, freedom, and artificiality. This motif extended from the lighting and interior setup to drinks and desserts, invoking a sense of alien futurity. Table settings included napkins (for the Ortolan ritual), feathers, bird dolls, and glassware shaped like birds. These theatrical elements were chosen for their symbolic and sensory resonance, blending natural motifs with synthetic speculation. All guests were instructed to place napkins over their heads during the main course, reviving the traditional Ortolan ritual in a new technological context.
Culinary Design and Symbolism: The three-course menu (see Figure 10) reflected the narrative arc of the drama and incorporated speculative and symbolic ingredients. The starter mimicked the natural diet of Ortolans, placing participants in the imagined perspective of the bird. The main course substituted Ortolan with quail for ethical, legal, and sensory alignment, allowing diners to experience a taste analogous to the original dish (Poon, 2014). The final dessert—a futuristic blue concoction—reinforced the speculative setting and closed the meal with a sense of surrealism. Real ingredients were disclosed in advance to respect dietary needs and ethical transparency.
Further details of the visual and morphological design process, including intermediary sketches and iterative explorations of wings, head structures, and textural elements, are provided in  Appendix B. These materials complement the performative account presented here by documenting the broader development of the fictional biohybrid creature.
3.4. Participants and Their Role Play
We recruited six guest participants using snowball sampling, all professionals working in creative, design, and cultural sectors. This sampling was intentional: creative professionals often serve as “lead users” with imaginative capacities and critical reflection skills to engage productively in speculative inquiry. The participants were: P1, a design lecturer at a world-renowned institution; P2, an award-winning poet; P3, an artist and gallery owner; P4, a foresight expert from a global tech innovation hub; P5, a professional UX designer; and P6, a filmmaker. The group was gender-balanced, with five participants in their 30s and one in his 70s. Participants were verified as meat eaters, and food allergies were checked in advance. One participant with dietary restrictions participated as a non-eater, taking on the role of the restaurant service manager.
@@ 3.5. Event Procedure
Figure 3. Procedure diagram.
A nicely illustrated flowchart.
@@
The host (lead artist) welcomed guests and introduced the speculative narrative. The chef (lead researcher) then explained the fictional menu and contextualized the key concept of edible animal-robot hybrids. Dramatic beats followed: a feathered drone representing the “synthetic ortolan” was flown across the room (Figure 4d), after which it was replaced by a 3D prototype passed among the guests (Figure 4e). The final edible course—a roasted quail representing the hybrid—was served under napkins, referencing the historic Ortolan tradition (Figure 4f). This sequence was designed to gradually increase sensory immersion, shifting participants from observational to embodied engagement, and culminating in reflective dialogue. The experience concluded with a dessert and a moderated post-dinner discussion. The event was video-recorded using a 360° omnidirectional camera, and the researchers took field notes throughout the session.
Different moments of the event that a group of people were participating.
Figure 4. Highlights from the dinner-in-the-drama: (a) Opening remarks; (b) Introducing key concepts; (c) Enjoying the show; (d) Biohybrid robot in flight; (e) Observing the biohybrid robot; (f) Serving biohybrid robots; (g) Eating biohybrid robots.